OpenStudy (errinss):

i've been graphing pretty much every x and y axis, but i still dont understand how to graph this can you help will fan and medal Graph f(x)=−0.25x+4

HanAkoSolo (jamierox4ev3r):

Here are a few key thing to look at **Slope: is it negative or positive? **y-intercept: this will tell you where on your graph the slope will begin

OpenStudy (errinss):

its negative

OpenStudy (triciaal):

it will be easier to express the decimal as a fraction follow the above

HanAkoSolo (jamierox4ev3r):

or rather, the y intercept will tell you where on the y axis you will begin your slope

OpenStudy (errinss):

im so confused

HanAkoSolo (jamierox4ev3r):

so the above user was saying that you should turn the slope into a fraction so -0.25x = -\(\Large\frac{1}{4}\)

HanAkoSolo (jamierox4ev3r):

**-\(\Large\frac{1}{4}\)x

HanAkoSolo (jamierox4ev3r):

remember, to find the slope, you want to find the rise over run. \(\Large\frac{rise}{run}\) --> the numerator would be your "rise." so if we have a numerator of -1, the line goes down one for every run. The run is 4.

HanAkoSolo (jamierox4ev3r):

I can illustrate this for you if you'd like, but is everything making a little more sense now?

OpenStudy (errinss):

yes please can you

HanAkoSolo (jamierox4ev3r):

sure. 

HanAkoSolo (jamierox4ev3r):

Does this illustration help?